### v1

# Language Repository for Long Video Understanding

Abstract
Language has become a prominent modality in computer vision with the rise of multi-modal LLMs. Despite supporting long context-lengths, their effectiveness in handling long-term information gradually declines with input length. This becomes critical, especially in applications such as long-form video understanding. In this paper, we introduce a Language Repository (LangRepo) for LLMs, that maintains concise and structured information as an interpretable (i.e., all-textual) representation. Our repository is updated iteratively based on multi-scale video chunks. We introduce write and read operations that focus on pruning redundancies in text, and extracting information at various temporal scales. The proposed framework is evaluated on zero-shot visual question-answering benchmarks including EgoSchema, NExT-QA, IntentQA and NExT-GQA, showing state-of-the-art performance at its scale. Our code is available at github.com/kkahatapitiya/LangRepo.
Keywords:
Large Language Models Long-form Video
1 Introduction
Video data is central to learning systems that can interact and reason about the world. Yet, they also associate with significant challenges such as increased compute requirements and redundant information, to name a few. This is especially critical in long-term videos. Even so, recent literature on video understanding have progressed so far, enabling reasoning capabilities in hours-long video streams [72, 27], in contrast to very-limited temporal spans (e.g. seconds or minutes) just a few years ago. Such progress is intriguing considering how complex the semantics become when temporal span is increased [68, 89]. Work on efficient spatio-temporal attention mechanisms [2, 4], memory management [79, 63], and large-language-models (LLMs) [76, 90, 72] have been key ingredients for such recent improvements.
LLMs, or more-specifically, vision-large-language-models (VLLMs) have been outperforming pure vision models in recent years in all facets, including image-based reasoning [49, 97, 43], grounding [38, 60], video understanding [76, 88, 90], and even robotics [91, 1, 46]. The sheer model scale and the vast pretraining data have enabled such frameworks to capture world knowledge and semantics, beyond what is possible with visual data only. Besides, the ability to process long context-lengths is also key, as it helps modeling long-term dependencies that are crucial for more-complex reasoning and interactions. However, recent studies show that despite the availability of such context-lengths, the effectiveness of models declines with longer input sequences [41]. This promotes the search for alternate representations that can compress input language data without losing meaningful information, essentially managing the context utilization of LLMs.
Moreover, the use of text (i.e., language) in modeling has shown numerous benefits such as rich semantics [77, 52, 31], ease of information sharing between different specialized-models [91] or modalities [49, 21], and interpretability [94, 69]. Among such, interpretability has a huge societal impact in the age of LLMs, to manage adversities such as bias [47, 18] and hallucinations [93, 12]. Simply put, it enables human observers to understand and monitor what really happens within models. Hence, interpretable representations have also been of interest to the community, in place of latent representations [79, 63].
Figure 1: Overview of our Language Repository (LangRepo): We propose an all-textual repository of visual information that updates iteratively, creating a multi-scale and interpretable representation. It extracts information from captions (generated by a VLLM [49, 96]) corresponding to video chunks. In write-to-repo, we prune and rephrase input descriptions, creating concise entries in the repository. In read-from-repo, such language descriptions (together with any optional metadata, e.g., timestamps) at multiple semantic-scales are summarized to generate outputs suited for video VQA. Here, rephrase and summarize are LLM-calls. We also compare LangRepo against state-of-the-art methods, showing strong performance at its scale.
Motivated by the above, we introduce Language Repository (LangRepo), an interpretable representation for LLMs that updates iteratively. It consumes input captions corresponding to video chunks , as shown in Fig. 1 (left). As LangRepo is all-textual, we rely on text-based operations to write and read information. The write operation (write-to-repo) prunes redundant text, creating concise descriptions that keep the context-utilization of LLMs in-check. Its iterative application with increasingly-longer chunks enables it to learn high-level semantics (e.g. long temporal dependencies). The read operation (read-from-repo) extracts such stored language information at various temporal scales, together with other optional metadata within the repository entries (e.g. timestamps). Altogether, our proposed framework is applied to long-term video reasoning tasks including visual question-answering (VQA) on EgoSchema [51], NExT-QA [80] and IntentQA [42], and visually-grounded VQA on NExT-GQA [81], showing strong performance at its scale, as given in Fig. 1 (right). Finally, we ablate our design decisions, providing insights on key components.
2 Related work
2.0.0.1 Long-video understanding:
Video models have progressed over the years, going from primitive recognition tasks [70, 36] to complex and fine-grained reasoning tasks [68, 89, 80, 22, 51] over long horizons. Both convolutional baselines [6, 17, 16] and transformer architectures [2, 4, 54] have explored research directions such as multi-scale representations [17, 14, 50], efficiency concerns associated with heavy spatio-temporal computations [13, 44], and handling redundant information within video inputs [8, 32]. More recently, long-video understanding has made a leap forward thanks to benchmark datasets [22, 51, 80] and model improvements [90, 92, 56], validating the importance of modeling complex interactions that happen over long periods of time. Still, the sub-par performance of SOTA models on such benchmarks suggests the room for improvement.
2.0.0.2 Long-context models:
Even before the age of LLMs, models based on convolutions [75, 58, 57, 32], recurrent blocks [24, 10, 26] or transformers [79, 63, 7] have exploited long-term dependencies, especially in the context of video understanding [75, 79] and reinforcement learning [7, 65]. With the rise of LLMs, scaling laws have revealed the importance of longer contexts even more [72, 61], and, thanks to the breakthroughs such as sparse processing [66, 15], caching [20, 37, 33], model-sharding [95, 9, 40], and efficient attention [11, 39], such long-context LLMs have become a reality. Even with very large context lengths, maintaining the effectiveness of reasoning over longer inputs is challenging [41, 85, 67]. This motivates us to think about concise representations that can better-utilize LLM context.
2.0.0.3 Compressing representations:
When handling heavy inputs, deep learning models have relied on compressed representations. It may come in the form of pruning [64, 5], latent memory [63, 23, 79], or external feature banks [78], to name a few. Despite the intuitive novelties and efficiency gains of such techniques, it is challenging to realize which information get preserved, and how semantically-meaningful they are post-compression. An interpretable representation that supports compression, if available, may shed light on such details.
2.0.0.4 Language as an interpretable modality:
More-recently, language has emerged as a dominant modality in computer vision due to its strong generalization capabilities [59, 28]. It can also act as a bridge between various domain-specific models [91], other modalities [49, 21], and even human instructions [71, 25], showing intriguing applications in domains such as chat agents (e.g. ChatGPT, Bard) and robot-learning [1, 46]. Since language is interpretable, it enables humans to interact with models more-naturally and make sense of model predictions.
Motivated by the above, we introduce an interpretable language representation that can (1) prune redundant information, and (2) extract multi-scale (or, high-level) semantics, enabling better context-utilization within LLMs.
3 Observations on long-range inputs
In this section, we investigate how LLMs perform with increasing inputs lengths (i.e., #tokens). Recent LLMs with very-large context lengths such as Gemini-Pro-1.5 [72] (1M tokens) or Claude-2.1 (200k tokens), can support extremely long input sequences. Yet, when feeding longer inputs, the reasoning capabilities (especially, long-term reasoning) of such models diminish. This behavior is also observed in concurrent work [41], and evident in benchmark results of state-of-the-art models [88, 90] (i.e., better performance with shorter inputs, or fewer video frames). To better investigate this in our setup, we evaluate VQA performance on standard long-term video understanding benchmarks while varying the input length (see Table 1). We consider frame/short-clip captions extracted using a VLLM at a baseline framerate (1$\times$) as inputs (introduced in [92]). We either subsample (0.5$\times$) or replicate (2$\times$) the captions, decreasing/increasing the input lengths of a question-answering LLM, namely, Mistral-7B [29] with 8k (or, theoretical 128k) context length. All inputs fit within the context, without any overflow. The observation from this study is consistent: even though the context length of the LLM is sufficient to process given inputs, the effectiveness of its predictions (shown by VQA performance) drops with longer inputs. This motivates us to introduces a concise language representation that preserves important details of long-range inputs, while pruning any redundant information.
Table 1: Observations on increasing input length: We evaluate the performance of a question-answering LLM (Mistral-7B [29]) at different input lengths, on multiple long-video reasoning benchmarks: EgoSchema [51], NExT-QA [80] and IntentQA [42]. Here, 1×1$\times$1 × corresponds to captions extracted by a VLLM at a default frame-rate (and, 0.5×0.5$\times$0.5 ×/2×2$\times$2 × corresponds to a compression/expansion of them by a factor of 2). We see that even though the context length of the LLM can fit all inputs without any overflow, the effectiveness of predictions (i.e., VQA performance) decreases with input length.
Dataset Captions per-video
0.5$\times$
1$\times$
2$\times$
EgoSchema (subset) [51]
49.8
48.8
46.8
NExT-QA [80]
48.2
48.2
46.9
IntentQA [42]
47.1
46.9
45.2
4 Language Repository
Figure 2: Detailed view of our Language Repository (LangRepo): Here we present the write and read operations within LangRepo. Given short-captions corresponding to video chunks, write-to-repo first prunes redundant captions within each chunk. The same process is iteratively applied on increasingly longer (or, higher-level) chunks— that are already within the repository— to generate multi-scale repository entries. Pruning consists of two stages: (1) grouping most similar captions based on embedding (eg: CLIP [59]) similarities between two subsets, and (2) rephrasing grouped captions with an LLM-call. The resulting LangRepo will include rephrased-captions and any optional metadata (e.g. #occurrences, timestamps). Next, read-from-repo generates concise descriptions for different semantic levels by summarizing the multi-scale language representation, which is also an LLM-call.
We present a Language Repository (LangRepo) that iteratively updates with multi-scale descriptions from video chunks. In contrast to external feature banks [78] or learnable latent memory representations [79, 63, 3], our proposal has a few key advantages: (1) it requires no training (i.e., zero-shot), and (2) it is compatible with both LLM-based processing and human interpretation, as it is fully-textual, i.e., it exists in language-space instead of a latent-space. LangRepo consists of two main operations: (1) information writing (write-to-repo), which prunes redundancies and iteratively updates language descriptions based on increasingly-longer video chunks, and (2) information reading (read-from-repo), which extracts preserved descriptions (with any optional metadata) in multiple temporal scales. We show a detailed view of these operations in Fig. 2, and further elaborate in the following subsections.
Consider a long video that is split in to $n$ non-overlapping chunks, denoted as $V=\{v_{i}\;|\;i=1,\cdots,n\}$. Assume that we already have frame or short-clip captions extracted by a VLLM (e.g. LLaVA [49]) corresponding to such chunks, denoted by $C=\{c_{i}\;|\;i=1,\cdots,n\}$. Here, each chunk may consist of $p$ such captions as in $c_{i}=\{c_{ij}\;|\;j=1,\cdots,p\}$. Altogether, $V$ is represented by $n\times p$ captions which we consider as inputs to our framework.
4.1 Writing to repository
[…]
In each subsequent iteration $k$, previous repo entries of iteration $k-1$ are re-combined into longer chunks and processed in the same way, generating information for higher semantic-levels.
[…]
Here, re-chunk($\cdot$) denotes the creation of longer chunks within the repository. More specifically, we simply concatenate (denoted by []delimited-[][$\cdot$][ ⋅ ]) all entries from previous iteration, and split them again into fewer number of chunks (hence, longer chunk size). Note that $i^{\prime}$ in the above equation is not the same as the previous chunk indexing $i$, as we may have different (usually, fewer) number of chunks in each subsequent iteration.
Each write operation involves two stages: (1) Grouping redundant text, and (2) Rephrasing, which are detailed below.
4.1.0.1 Grouping redundant text:
Given textual descriptions of a video chunk (i.e., captions in the first write iteration, or previous repo descriptions in subsequent iterations), we plan to identify most-similar ones and merge them as a single description. Without loss of generality, let us consider the first write iteration, for which the input is in the form of $c_{i}=\{c_{ij}\;|\;j=1,\cdots,p\}$. Inspired by [5], we first split the captions of each chunk into two sets, namely, source (src) captions $c_{\text{src},i}$ and destination (dst) captions $c_{\text{dst},i}$. Let us drop the chunk index $i$ for brevity. Here, dst captions $c_{\text{dst}}$ are sampled uniformly distributed across the temporal span of a chunk, while all the rest are considered as src captions $c_{\text{src}}$ (see Fig. 2).
$c_{\text{dst}},\;c_{\text{src}}$ $=\texttt{split}(c).$ (3)
Here, we usually have fewer dst captions (i.e., $|c_{\text{dst}}|<|c_{\text{src}}|$). Next, we embed all captions using a text-encoder (e.g. CLIP [59]), and compute the cosine similarity of each pair between src-dst sets to find most-similar matches.
$\text{sim}_{\text{src-dst}}$ $=\texttt{sim}(\texttt{encode}(c_{\text{src}}),\;\texttt{encode}(c_{\text{dst}})).$ (4)
Based on the similarity matrix above ($\text{sim}_{\text{src-dst}}$), we then prune the highest $x\%$ similarities by grouping such source captions with their corresponding destination matches, forming a set of grouped descriptions $c_{\text{grp}}$ for the given chunk. Refer to the color-coded captions after ‘Group’ in Fig. 2.
$c_{\text{grp}}$ $=\texttt{group}(c_{\text{dst}},\;c_{\text{src}},\;\text{sim}_{\text{src-dst}},\;x).$ (5)
Here $x\%$ is a pre-defined hyperparameter (e.g. 50%). Finally, such grouped descriptions go through a rephrasing operation prior to entering the repository.
Figure 3: LLM prompt templates in LangRepo: Here, we show the zero-shot prompt templates used for rephrasing ($\texttt{template}_{\texttt{reph}}$) and summarizing ($\texttt{template}_{\texttt{sum}}$) operations. Rephrase prompt needs a list of grouped captions as input, while its output adheres to more-strict requirements (e.g. same order, same number of list items) needed for correct parsing. Summarize prompt takes in each repository entry and generates a more-flexible (i.e., open-ended) output, while also conditioning on the question.
4.1.0.2 Rephrasing:
Grouped captions $c_{\text{grp}}$ of each chunk are rephrased via an LLM-call. This allows redundant information within each group to be dropped, while generating a concise and coherent description. We first form a list of grouped captions, where each list item corresponds to a single group (i.e., a dst caption and any one or more src captions matched to it), and feed it to the LLM, wrapped in a rephrasing-template ($\texttt{template}_{\texttt{reph}}$) as shown in Fig. 3 (top).
${c_{\text{reph}}}$ $=\texttt{rephrase}(\texttt{template}_{\texttt{reph}}(c_{\text{grp}},\;l)).$ (6)
Here, $l$ denotes the number of grouped captions in the given chunk. The LLM output (${c_{\text{reph}}}$) is restricted to be a list in the same order with the same number of items, where each item is a single concise sentence. Finally, such rephrased descriptions together with other metadata such as timestamps ($t$) and number of occurrences ($o$) are written in the repository.
$r^{1}$ =$=\{(c_{\text{reph},j},\;t_{j},\;o_{j})\;|\;j=1,\;\cdots,\;p^{\prime}\}.$ (7)
Note that here $p^{\prime}<p$ as we have grouped and rephrased $x\%$ of most-similar captions. Alongside each description in a repository entry, $t$ maintains a growing list of timestamps corresponding to its founding captions, whereas the occurrences counter ($o$) keeps track of the number of such captions. A qualitative example of a repository entry is given in Fig. 4.
In subsequent iterations, the same operations apply when writing multi-scale entries. The only difference is the change in input, which now constitutes of previous repo entries re-combined into high-level chunks (i.e., $c\rightarrow\texttt{re-chunk}([\cdots,\;r^{k-1},\;\cdots])$ in Eq. 3). Each new iteration generates information corresponding to a higher semantic-level (i.e., going from short-range to long-range dependencies), forming our multi-scale language representation.
Figure 4: A qualitative example of a LangRepo entry: Given a video chunk, redundant captions are first grouped together during pruning operation. During rephrasing, such groups are more-concisely written to the repository, along with additional metadata. Other non-redundant captions are written directly. This process is continued iteratively with increasingly-longer chunks, creating multi-scale repository entries (refer Fig. A.1 for a more-detailed view). Finally, such descriptions from various temporal scales are read to generate the output.
4.2 Reading from repository
As we make a single VQA prediction for a given long video— instead of making predictions every chunk— our read operation (read-from-repo) is applied only after fully-forming the multi-scale repository (i.e., after all write iterations). The repo entries after $K$ write iterations can be denoted as $\{r^{k}\;|\;k=1,\cdots,K\}$ where each scale ($r^{k}$) may consist of multiple entries $\{\cdots,\;r^{k}_{i-1},\;r^{k}_{i},\;r^{k}_{i+1},\;\cdots\}$. When reading, we generate summaries for each entry in the repo separately, allowing it to focus on varying temporal spans. More specifically, each entry goes through a summarizing-template ($\texttt{template}_{\texttt{sum}}$) as shown in Fig. 3 (bottom), and the resulting prompt is fed to the LLM.
[…]
Here, $d^{k}_{i}$ corresponds to the output description of each entry $i$ in the repository, at the respective scale $k$. Optionally, we can make use of additional metadata such as timestamps and #occurrences, by prompting the read operation with descriptions of repo entries formatted as “[timestamps] description ($\times$#occurrences)” (see Fig. 4). Finally, we concatenate all output descriptions and prompt the LLM again to generate the answer prediction.
ans $=\texttt{vqa}([\cdots,d^{k}_{i},\cdots]).$ (9)
4.3 Design decisions
4.3.0.1 Similarity-based pruning:
We notice that the short captions generated by the VLLM can be highly-redundant, as it has a limited temporal span. Such excess details can adversely affect the performance (see Table 1), while also wasting the LLM context. This motivates us to prune redundancies. We consider prompting the LLM directly to identify and rephrase redundant information. However, the outputs in this setup can be noisy and lack of any structure that is useful for parsing. In other words, although redundancies get pruned, there is limited controllability and inability of identifying what gets pruned. Hence, we decide to delegate the function of identifying redundancies to a separate module: a similarity-based grouping with the help of text embeddings. This gives more control on what to prune and how much to prune, while generating outputs that can be parsed to extract other useful metadata (e.g. timestamps).
4.3.0.2 Processing videos as chunks:
Our decision to consume longer videos as chunks is motivated by prior work [79, 63]. It allows us to not lose short-term details, while also keeping track of long-term dependencies via multi-scale processing. Additionally, although not explored in the scope of this paper, such a setup integrates well with temporally-fine-grained prediction tasks, where an LLM needs to make multiple predictions over time.
4.3.0.3 Choice of metadata:
To avoid the loss of important details during pruning, we maintain additional metadata in our LangRepo. Since captions across time can be grouped together in a single repo description, we save their timestamps as a separate field. This can help with temporal reasoning questions. We also update an occurrence counter, which shows the number of captions grouped within a single description. This can act as a weight, to help in cases such as counting or identifying repetitive events.
4.3.0.4 All-textual repository:
Instead of being a latent representation [79, 63, 3], our LangRepo is all-textual. This promotes interpretability for human observers, while also being a more-natural form of structure for LLM-based processing. Additionally, our implementation can be formulated to be zero-shot, without requiring any training or finetuning.
4.3.0.5 Classifier for close-ended VQA:
The standard multiple-choice question-answering setup considers a generative classifier. Meaning, an LLM is prompted to generate the correct answer option among multiple-choices, directly as next-token prediction. Another approach used in NLP literature is the log-likelihood based classification. Here, the LLM is prompted separately for each of the multiple choices with a template such as “Question: Answer-option”. The choice that maximises the log-likelihood of predicted tokens (i.e., tokens corresponding to Answer-option) is selected as the correct answer. This is a more-natural setup for close-ended VQA since it avoids hallucination. Among these classifiers, we find the latter to be better-performing. Yet, it is more-sensitive to the prompt template. We direct the reader to supplementary A.1 for more details.
5 Experiments
5.1 Implementation details
In our experiments, we rely on frame or short-clip captions pre-extracted using VLLMs, as given in [92]. For EgoSchema [51], LaViLa [96] is used to generate captions at 1fps. For NExT-QA [80] data used in other benchmarks, LLaVA-1.5 [49] generates captions at 0.5fps. By default, we use either Mistral-7B [29] (w/ 7B parameters) or Mixtral-8$\times$7B [30] (w/ 12B active parameters) as our LLMs. As the text encoder in similarity-based pruning, we use CLIP-L/14 [59]. Note that all the models used in our framework are open-source and within a reasonable model-scale, promoting further research in academic settings. We do zero-shot inference on all datasets without any finetuning, evaluating the performance on long-term video VQA benchmarks.
5.2 Dataset details
5.2.0.1 EgoSchema:
EgoSchema [51] is a long-video VQA dataset derived from Ego4D [22], by semi-automatically generating and verifying mutliple-choice questions. Each video is 3 minutes long, and the questions are filtered to have a long temporal certificate (i.e., at least 100 seconds of a video is required to answer the question). The public validation subset consists of 500 videos, each with 5 answer-choices and the correct ground-truth answer. The fullset of 5K videos is hosted as a Kaggle challenge. There is no training split for this dataset, motivating zero-shot evaluation. In our experiments, we report the performance for both subset and fullset.
5.2.0.2 NExT-QA:
NExT-QA [80] is a popular VQA dataset with videos up to 2 minutes long, at an average of 44 seconds. It consists of 52k open-ended questions and 48k close-ended questions (i.e., multiple-choice with 5 answer options) in total. The questions are further classified into temporal, causal, or descriptive categories, to enable evaluation of different reasoning capabilities of models. We consider zero-shot evaluation on validation set of 5k multiple-choice questions, and report the performance on each category.
5.2.0.3 IntentQA:
IntentQA [42] is based on the same NExT-QA videos (specifically, the ones for temporal and causal reasoning), yet re-purposed focusing on intent-related questions. A new set of 16k multiple-choice questions are provided for 4.3k videos, each with 5 answer choices. These questions are further classified as why?, how? or before/after, based on the appearance of the same keywords in the question. In our evaluation, we focus on zero-shot setting on test set of 2k questions, reporting performance for each category.
5.2.0.4 NExT-GQA:
NExT-GQA [81] is a visually-grounded VQA dataset, which provides additional 10.5K temporal grounding annotations on top of original NExT-QA [80] data. The test split contains 5.5k question-answer pairs, together with grounding labels that convey the temporal span needed to answer the question. Here, we focus on zero-shot inference similar to [92], and report multiple metrics including Intersection-over-Prediction (IoP) that measures the overlap w.r.t. the predicted window, Intersection-over-Union (IoU) that measures the overlap w.r.t. the union of ground-truth and predicted windows, and Acc@GQA that measures the accuracy of correctly-grounded predictions.
5.3 Main results
Table 2: Results on EgoSchema [51]: We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). We focus on the zero-shot setting. LangRepo shows a strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
Model Params  Subset (%)  Fullset (%)
finetuned
MC-ViT-L [3] 424M 62.6 44.4
ImageViT [56] 1B 40.8 30.9
ShortViViT [56] 1B 47.9 31.0
LongViViT [56] 1B 56.8 33.3
zero-shot
VIOLET[19] 198M - 19.9
InternVideo [76] 478M - 32.1
FrozenBiLM [86] 890M - 26.9
[…]
mPLUG-Owl [88] 7B - 31.1
LLoVi [92] 7B 50.8 33.5
Vamos [74] 13B - 36.7
[…]
LangRepo (ours) 7B 60.8 38.9
LangRepo (ours) 12B 66.2 41.2
5.3.0.1 EgoSchema:
In Table 2, we present the VQA performance of LangRepo on standard EgoSchema [51] splits, comparing with other state-of-the-art frameworks. Here, we focus on zero-shot evaluation, yet also report finetuned setting (i.e., any downstream-data-specific training) for completeness. We consider Mistral-7B [29] and Mixtral-8$\times$7B [30] as the choice of LLMs in our setup, both with reasonable model scales (7B and 12B active parameters, respectively). We de-emphasize models with significantly-higher #parameters (e.g. 175B GPT-3.5, or 1.8T GPT-4 variants), and the ones that use video-caption pretraining. LangRepo shows significantly-better performance compared to other methods at a similar scale, validating its effectiveness in long-video reasoning. We achieve $+10.1\%$ on fullset compared to mPLUG-Owl [88], $+15.4\%$ on subset and $+7.7\%$ on fullset compared to LLoVi [92] (w/ Mistral-7B [29]), and $+4.5\%$ on fullset compared to Vamos [74] (w/ Llama2-13B [73]).
Table 3: Results on NExT-QA [80]: We compare LangRepo against state-of-the-art methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
 Model Params  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
CoVGT [84] 149M 58.8 57.4 69.3 60.0
SeViTFiD [34] 215M - - - 60.6
HiTeA [87] 297M 62.4 58.3 75.6 63.1
MC-ViT-L [3] 424M - - - 65.0
InternVideo [76] 478M 62.5 58.5 75.8 63.2
BLIP-2 [43] 4B 70.1 65.2 80.1 70.1
[…]
LLama-VQA [35] 7B 72.7 69.2 75.8 72.0
Vamos [74] 7B 72.6 69.6 78.0 72.5
zero-shot
VFC [53] 164M 45.4 51.6 64.1 51.5
InternVideo [76] 478M 43.4 48.0 65.1 49.1
[…]
Mistral [29] 7B 51.0 48.1 57.4 51.1
LLoVi [92] 7B 55.6 47.9 63.2 54.3
[…]
LangRepo (ours) 7B 57.8 45.7 61.9 54.6
LangRepo (ours) 12B 64.4 51.4 69.1 60.9
5.3.0.2 NExT-QA:
In Table 3, we report the performance of LangRepo on standard NExT-QA [80] validation splits (Causal, Temporal and Descriptive) and the full validation set. On zero-shot evaluation, our framework outperforms other methods consistently. Compared to smaller models, we gain $+11.8\%$ over InternVideo [76] and $+9.4\%$ over VFC [53]. Compared to models of similar scale, we gain $+9.8\%$ over baseline Mistral-7B [29] and $+6.6\%$ over LLoVi (w/ Mistral-7B) [92]. We de-emphasize LLoVi variant with GPT-4 due to its whopping scale of 1.8T parameters. We also de-emphasize SeViLA [90] which is pretrained with video-captions, whereas we rely on LLaVA-1.5 [48] captions that has not seen any video pretraining. Finally, we observe that LangRepo outperforms competition on semantic splits showing the generalization of our language representation.
Table 4: Results on IntentQA [42]: We compare LangRepo against state-of-the-art methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
Model Params  Why? (%)  How? (%)  Before/After (%)  All (%)
finetuned
HQGA [82] 46M 48.2 54.3 41.7 47.7
VGT [83] 511M 51.4 56.0 47.6 51.3
Vamos [74] 7B 69.5 70.2 65.0 68.5
[…]
zero-shot
[…]
Mistral[29] 7B 52.7 55.4 41.5 50.4
LLoVi [92] 7B 57.9 55.4 42.3 53.6
[…]
LangRepo (ours) 7B 56.9 60.2 42.1 53.8
LangRepo (ours) 12B 62.8 62.4 47.8 59.1
5.3.0.3 Intent-QA:
In Table 4, we evaluate our zero-shot framework against other state-of-the-art models on Intent-QA [42] test splits (Why?, How? and Before/After) and the full test set. LangRepo outperform comparable models with similar scale consistently, showing gains of $+8.7\%$ over baseline Mistral-7B [29] and $+5.5\%$ over LLoVi [92] (w/ Mistral-7B). Again, we de-emphasize significantly larger models (LLoVi w/ GPT-4) and models pretrained with video-captions (SeViLA [90]), whereas ours rely on LLaVA-1.5 [48] frame-level captions.
Table 5: Results on NExT-GQA [81]: We compare LangRepo against state-of-the-art methods on NExT-GQA test set. We focus on the zero-shot setting. Our method shows strong performance at its scale. Models with significantly higher #parameters or video-caption ${}^{\dagger}$ are de-emphasized.
 Model Params mIoP IoP@0.5 mIoU IoU@0.5 Acc@GQA
weakly-supervised
IGV [45] 110M 21.4 18.9 14.0 9.6 10.2
Temp[CLIP] [59, 81] 130M 25.7 25.5 12.1 8.9 16.0
FrozenBiLM [86, 81] 1B 24.2 23.7 9.6 6.1 17.5
[…]
zero-shot
Mistral [29] 7B 20.4 20.2 8.7 5.9 9.2
LLoVi [92] 7B 20.7 20.5 8.7 6.0 11.2
[…]
LangRepo (ours) 7B 20.3 20.0 8.7 6.0 11.2
LangRepo (ours) 12B 31.3 28.7 18.5 12.2 17.1
5.3.0.4 NExT-GQA:
In Table 5, we compare the performance of LangRepo with state-of-the-art models on NExT-GQA. Our method achieves a strong performance at its scale, outperforming baseline Mistral-7B [29] by $+7.9\%$ and LLoVi [92] (w/ Mistral-7B) by $+5.9\%$ on Acc@GQA metric. Despite being zero-shot, it is also competitive with weakly-supervised baselines. Here, we de-emphasize significantly-larger models (LLoVi w/ GPT-4) and models pretrained with video-captions (SeViLA [90]).
5.4 Ablation study
Table 6: Ablations on EgoSchema [51]: We evaluate different design decisions of our framework on EgoSchema 500-video subset for zero-shot video VQA.
LLM   Scale Acc. (%)
Llama2 [73] 13B 43.0
Mistral [29] 7B 50.8
Text encoder Acc. (%)
Sentence-T5-XL [62]
56.4
CLIP-L/14 [59]
57.8
VQA classifier       Acc. (%)
Genearative 57.8
Log-likelihood 60.8
#Iterations  #Chunks  Scales-read  Acc. (%)
1 [2] [1st] 57.0
1 [4] [1st] 60.8
2 [3,2] [2nd] 55.6
2 [3,2] [1st, 2nd] 60.6
3 [4,3,2] [3rd] 58.4
3 [4,3,2] [2nd, 3rd] 59.4
3 [4,3,2] [1st, 2nd, 3rd] 61.2
Model
Acc. (%)
LangRepo (ours)
60.8
+\;\;++ timestamps
60.4
+\;\;++ occurrences
61.4
+\;\;++ timestamps $+$ occurrences
58.2
Streaming setup Acc. (%)
LLoVi [92]
50.8
Chunk-based LLoVi
57.8
LangRepo (ours)
60.8
Model
0.5$\times$
1$\times$
2$\times$
Mistral-7B [29]
49.8
48.8
46.8
LLoVi [92]
57.2
55.4
53.6
LangRepo (ours)
56.4
57.8
56.4
5.4.0.1 Choice of backbone LLM, text encoder and classifier:
We consider different open-source LLM-backbones in our setup, namely, LLama2 [73] and Mistral [29]. We ablate the choice of LLM within LLoVi [92] framework, as shown in Table LABEL:tab:ablation:llm_choice. We observe that Mistral-7B is significantly better at video reasoning compared to LLama2-13B. Next, we consider different text encoders to embed our text descriptions prior to pruning. Namely, we use either CLIP-L/14 [59] or Sentence-T5-XL [62]. As presented in Table LABEL:tab:ablation:text_encoder. Surprisingly, CLIP outperforms Sentence-T5 that is trained with a sentence-level objective (which is expected to better align with our caption-similarity computation). Finally, we evaluate different classifiers used for close-ended (i.e., multiple-choice question) VQA setups (see Table LABEL:tab:ablation:vqa_classifier). Even though the generative classifier is commonly-used in LLM literature, we observe that the log-likelihood classifier performs better. Such performance is also intuitive as the latter constrains predictions within the given answer choices (hence, no hallucination). Based on this observation, we adopt the log-likelihood classifier in our main experiments.
5.4.0.2 Repository setup and metadata:
In the formulation of LangRepo we ablate different hyperparameter settings related to the number of repo-updates (#iterations), the number of video chunks in each iteration (#chunks), and multiple temporal-scales considered when reading data in repository. In Table LABEL:tab:ablation:memory_setup, we make two observations: (1) more update iterations with finer chunks (higher #chunks per iteration) can preserve more-useful information, and (2) reading information in multiple temporal-scales is consistently better. Moreover, we consider optional metadata that can help preserve information that may get lost when pruning (e.g. temporal ordering, or repetitive captions), namely, timestamps and #occurrences (i.e., the number of captions grouped within each repo description). We see in Table LABEL:tab:ablation:metadata that #occurrences help weigh each description when summarizing, resulting in better performance. However, timestamps do not provide meaningful improvement in our setup, in the context of EgoSchema VQA.
5.4.0.3 Input format and length:
We also consider different ways of consuming long video data, either as a whole or as chunks (see Table LABEL:tab:ablation:streaming_setup). Among these options, processing as chunks enables preserving more fine-grained details in LLM outputs. Our repository setup provides further improvement showing its effectiveness over the baseline with the same chunk-based processing. Finally, we re-visit the experiment on how the input length affects the effectiveness of LLMs, presented in Table 1. In Table LABEL:tab:ablation:input_length, we show that LangRepo provide more-stable performance with increasing input lengths, in contrast to baselines, validating how it addresses this shortcoming.
6 Conclusion
In this paper, we introduced a Language Repository (LangRepo), which reads and writes textual information corresponding to video chunks, as a concise, multi-scale and interpretable language representation, together with additional metadata. Both our write-to-repo and read-from-repo operations are text-based and implemented as calls to a backbone LLM. Our empirical results show the superior performance of LangRepo on multiple long-video reasoning benchmarks at its respective scale, while also being (1) less-prone to performance drops due to increasing input lengths, and (2) interpretable, enabling easier human intervention if and when needed.
6.0.0.1 Acknowledgements:
This work was supported by the SUNY Empire Innovation Program (EIP). We thank the Robotics Lab at SBU for helpful discussions, and especially Xiang Li, for managing the lab servers and promptly responding to all the infrastructure-related issues.
A Appendix
A.1 Prompting for VQA
As the evaluation setup, we consider multiple-choice visual question-answering (VQA) on long videos. Given the close-ended answer formulation, we can consider two different classifiers to make the prediction: (1) a Generative classifier, which directly generates the answer choice, or (2) a Log-likelihood classifier, which select the most-probable choice based on the joint-probability of tokens in each answer option given the description and the question. As we discussed in Sec. 4.3, the latter generally performs better, as it is less-prone to hallucinations (i.e., prediction is explicitly constrained to answer choices). However, it is also sensitive to the prompts we use. Hence, we include a discussion on prompting in the following subsections.
A.1.0.1 Generative classifier:
Here, we direcly prompt the LLM to generate the correct answer, conditioned on the descriptions generated by LangRepo, the question and the answer options (inspired by [92]). To make sure that the output can be parsed, we provide additional guiding instructions and any syntax specific to the LLM (Mistral [29]). This also discourages any hallucinations. On all benchmarks, we use the common prompt given below.
[…]
A.1.0.2 Log-likelihood classifier:
[…]
A.2 Qualitative examples of repository entries
We present qualitative examples from EgoSchema [51] dataset to better clarify the operations in LangRepo. In Fig. 4, we show the format of repository entries. Here, non-redundant captions from the input get directly written to the repo. In contrast, any redundant captions— grouped based on similarity— get rephrased as concise descriptions (1 per-group). Each repository description may come with additional metadata such as timestamps and #occurrences to avoid the loss of meaningful information due to pruning. In Fig. A.1, we further elaborate on multiple scales within the repository, which are generated by iteratively processing increasingly-longer chunks (created by re-chunk operation). During reading, we can decide to summarize information at various temporal scales to generate output descriptions useful for VQA.
Figure A.1: A qualitative example of iterative writing and multi-scale reading in LangRepo: Here, we present an example with 2-scales, given captions of a 180s long video. In scale-1, we consider 3 chunks of 60s each, and in scale-2, we re-chunk them into 2 chunks of 90s each. We only show the redundant captions that go through pruning, and also, omit any metadata (e.g. timestamps) within the repository. In each scale, captions grouped based on similarity get rephrased concisely. To generate inputs of the subsequent scale, we simply order previous repository descriptions in time, and split (i.e., re-chunk) into fewer (and, longer) chunks. When reading, each entry in each scale is summarized separately to create output descriptions of various temporal spans. In general, we always consider the last-scale descriptions to be mandatory, but any prior-scale to be optional. Yet, we observe multiple scales to be beneficial (see Table LABEL:tab:ablation:memory_setup). Best-viewed with zoom-in.